# SSR-Speech: Towards Stable, Safe and Robust Zero-shot Text-based Speech Editing and Synthesis

### v1

@@ Abstract
In this paper, we introduce SSR-Speech, a neural codec autoregressive model designed for stable, safe, and robust zero-shot text-based speech editing and text-to-speech synthesis. SSR-Speech is built on a Transformer decoder and incorporates classifier-free guidance to enhance the stability of the generation process. A watermark Encodec is proposed to embed frame-level watermarks into the edited regions of the speech so that which parts were edited can be detected. In addition, the waveform reconstruction leverages the original unedited speech segments, providing superior recovery compared to the Encodec model. Our approach achieves the state-of-the-art performance in the RealEdit speech editing task and the LibriTTS text-to-speech task, surpassing previous methods. Furthermore, SSR-Speech excels in multi-span speech editing and also demonstrates remarkable robustness to background sounds. Source code [^1] https://github.com/WangHelin1997/SSR-Speech and demos [^2] https://wanghelin1997.github.io/SSR-Speech-Demo/ are released.
@@ I Introduction
Nowadays, zero-shot text-based speech generation [1, 2, 3, 4] has garnered significant attention in the speech community, particularly in areas like speech editing (SE) and text-to-speech (TTS) synthesis. Given an unseen speaker during training, zero-shot SE focuses on modifying specific words or phrases within an utterance to align with a target transcript while preserving the unchanged portions of the original speech, and zero-shot TTS is concerned with generating the whole speech following a target transcript. Recently proposed approaches based on large-scale speech data have significantly streamlined speech generation systems. Non-autoregressive (NAR) models, such as SoundStorm [5], FluentSpeech [6], NaturalSpeech 3 [7], and VoiceBox [8], have been proposed for their high inference speed and stability. However, they face challenges due to their reliance on phoneme-acoustic alignment and the complexity of the training process [9]. In contrast, language model (LM) based autoregressive (AR) models, such as VALL-E [10], UniAudio [11], and VoiceCraft [12], simplify the training process but are hindered by slow and unstable inference. For the SE task, existing methods struggle with handling multiple spans, speech with background noise or music, and preserving the unchanged portions effectively [13, 14, 15]. In addition, as these models can easily clone a human voice, AI safety becomes a potential concern [16, 17, 18].
In this work, we focus on AR models for zero-shot text-based SE and TTS, and proposed a novel Transformer-based AR model called SSR-Speech. The main contributions of this paper are summarized as follows:
(i) SSR-Speech leads to stable inference. Previous AR models may generate the long silence and scratching sound during generation, which produce unnatural sounding speech. The inference-only classifier-free guidance is applied to enhance the stability of the generation process.
(ii) The generated speech by SSR-Speech contains frame-level watermarks, which provides information whether the audio has been produced by SSR-Speech and which part of the audio has been edited or synthesised. To achieve this, a watermark Encodec model is proposed to introduce frame-level watermarks while reconstructing the waveform.
(iii) SSR-Speech is robust to multi-span editing and background sounds. The training pipeline of SSR-Speech includes single-span and multi-span editing, and editing any parts of the speech, so that there is no gap between training and inference for insertion, deletion and substitution. In addition, the watermark encodec leverages the original unedited speech segments for the waveform reconstruction, which provides better recovery compared to the Encodec model, especially for speech with background noise or music.
@@ II-B Inference
For the SE task, we compare the original transcript with the target transcript to identify the words that need to be masked. Using word-level forced alignment [^3] https://github.com/MontrealCorpusTools/Montreal-Forced-Aligner of the original transcript, we locate the corresponding masked spans of audio tokens. The phoneme tokens from the target transcript and the unmasked audio tokens are then concatenated and fed into the SSR-Speech model to autoregressively predict new audio tokens. Similar to [12], when editing speech, the neighboring words surrounding the span to be edited also need to be slightly adjusted to accurately model co-articulation effects. Thus, we introduce a small margin hyperparameter $\alpha$, extending the length of the masked span by $\alpha$ on both the left and right sides.
@@ IV-B Setups
Following [12], both the Encodec and Watermark Encodec models use 4 RVQ codebooks, each with a vocabulary size of 2048. They operate with a stride of 320 samples, resulting in a codec framerate of 50Hz for audio recorded at a 16kHz sampling rate. The base dimension is 64, doubling at each of the 5 convolutional layers in the encoder. The number of spans to be masked, denoted as $P$, is uniformly sampled between 1 and 3. The maximum masking length is set to $90\%$ of the original audio length. During training, we apply a probability of 0.5 to enhance TTS training. Text transcripts are phonemized using an IPA phoneset toolkit [^5] https://github.com/bootphon/phonemizer [27].
@@
For inference, we use nucleus sampling [29] with $p=0.8$ and a temperature of 1. The extended masked span $\alpha$ is set to $0.12$ seconds. Based on initial experiments, we determined the optimal value for the hyperparameter $\gamma$ to be $1.5$.
@@ IV-E Results
Table I presents the results of the speech editing evaluations on RealEdit. SSR-Speech outperforms the baselines across all metrics. From the ablation studies, we observed that CFG significantly contributes to the performance improvement, effectively resolving the long silence issue in the AR model. From the MOSNet-M results, this advantage is especially pronounced in multi-span editing. Comparing with VoiceCraft that would run inference for 10 times using different margin parameters, our proposed SSR-Speech is able to inference once and obtain a stable result.
Context-aware decoding (CD) also enhances performance, particularly in speech with background sounds, as the unedited spans provide additional context for the model. Therefore, SSR-Speech with CD gets much better MOSNet-N results than the others. While watermarking (WM) does not impact performance, it adds frame-level watermarks to the synthesized audio, increasing the model’s security.

### v2

@@ Abstract
In this paper, we introduce SSR-Speech, a neural codec autoregressive model designed for stable, safe, and robust zero-shot text-based speech editing and text-to-speech synthesis. SSR-Speech is built on a Transformer decoder and incorporates classifier-free guidance to enhance the stability of the generation process. A watermark Encodec is proposed to embed frame-level watermarks into the edited regions of the speech so that which parts were edited can be detected. In addition, the waveform reconstruction leverages the original unedited speech segments, providing superior recovery compared to the Encodec model. Our approach achieves state-of-the-art performance in the RealEdit speech editing task and the LibriTTS text-to-speech task, surpassing previous methods. Furthermore, SSR-Speech excels in multi-span speech editing and also demonstrates remarkable robustness to background sounds. The source code [^1] https://github.com/WangHelin1997/SSR-Speech and demos [^2] https://wanghelin1997.github.io/SSR-Speech-Demo/ are released.
@@ I Introduction
Nowadays, zero-shot text-based speech generation [1, 2, 3, 4] has garnered significant attention in the speech community, particularly in areas such as speech editing (SE) and text-to-speech (TTS) synthesis. Given an unseen speaker during training, zero-shot SE focuses on modifying specific words or phrases within an utterance to align with a target transcript while preserving the unchanged portions of the original speech, and zero-shot TTS is concerned with generating the whole speech following a target transcript. Recently proposed approaches based on large-scale speech data have significantly streamlined speech generation systems. Non-autoregressive (NAR) models, such as SoundStorm [5], FluentSpeech [6], NaturalSpeech 3 [7], and VoiceBox [8], have been proposed for their high inference speed and stability. However, they face challenges due to their reliance on phoneme-acoustic alignment and the complexity of the training process [9]. In contrast, language model (LM) based autoregressive (AR) models, such as VALL-E [10], UniAudio [11], and VoiceCraft [12], simplify the training process, but are hindered by slow and unstable inference. For the SE task, existing methods struggle to handle multiple spans, speech with background noise or music, and preserving the unchanged portions effectively [13, 14, 15]. In addition, since these models can easily clone a human voice, AI safety becomes a potential concern [16, 17, 18].
In this work, we focus on AR models for zero-shot text-based SE and TTS, and we proposed a novel Transformer-based AR model called SSR-Speech. The main contributions of this paper are summarized as follows:
(i) SSR-Speech leads to stable inference. Previous AR models may generate long silence and scratching noise during generation, which produce unnatural sounding speech. Inference-only classifier-free guidance is applied to enhance the stability of the generation process.
(ii) The generated speech by SSR-Speech contains frame-level watermarks, which provides information about whether the audio has been produced by SSR-Speech and which part of the audio has been edited or synthesized. To achieve this, a watermark Encodec model is proposed to introduce frame-level watermarks while reconstructing the waveform.
(iii) SSR-Speech is robust to multi-span editing and background sounds. The training pipeline of SSR-Speech includes single-span and multi-span editing, and editing any parts of the speech, so that there is no gap between training and inference for insertion, deletion, and substitution. In addition, the watermark encodec leverages the original unedited speech segments for the waveform reconstruction, which provides better recovery compared to the Encodec model, especially for speech with background noise or music.
@@ II-B Inference
For the SE task, we compare the original transcript with the target transcript to identify the words that need to be masked. Using word-level forced alignment [^3] https://github.com/m-bain/whisperX of the original transcript, we locate the corresponding masked spans of audio tokens. The phoneme tokens from the target transcript and the unmasked audio tokens are then concatenated and fed into the SSR-Speech model to autoregressively predict new audio tokens. Similar to [12], when editing speech, the neighboring words surrounding the span to be edited also need to be slightly adjusted to accurately model co-articulation effects. Thus, we introduce a small margin hyperparameter $\alpha$, extending the length of the masked span by $\alpha$ on both the left and right sides.
@@
Furthermore, we observed that CFG often generates speech at an accelerated pace due to the excessive removal of silence tokens during processing. To address this issue, we propose to utilize CFG with a stride of $\beta$ during inference, where $\beta$ serves as a hyperparameter.
1
2Function Inference-only CFG:
3      $\mathcal{M}$, $A^{\prime}$, $Y$, $Y^{\prime}$, $\beta$, $\gamma$, $[sog]$, $[eog]$
4
Input : Model $\mathcal{M}$, initial audio sequence $A^{\prime}$, target phoneme sequence $Y$, random phoneme sequence $Y^{\prime}$, stride $\beta$, guidance scale $\gamma$, start token $[sog]$, end token $[eog]$
Output : Generated sequence $S$
5 $S\leftarrow\textnormal{{concat(}}\textnormal{\emph{$A^{\prime}$, $[sog]$}}\textnormal{{)}}$;
6
[…]
8
9while True do
10       if $t\mod\beta=0$ then
11             $s_{t}\sim\gamma\mathbb{P}_{\theta}(S\mid Y)+(1-\gamma)\mathbb{P}_{\theta}(S\mid Y^{\prime})$;
12
13       end if
14      else
15             $s_{t}\sim\mathbb{P}_{\theta}(S\mid Y)$;
16
17       end if
18      $S\leftarrow\textnormal{{concat(}}\textnormal{\emph{$S$, $s_{t}$}}\textnormal{{)}}$;
19
20      if $s_{t}==[eog]$ then
21             break;
22
23       end if
24
25      $t=t+1$
26 end while
27
28return $S$;
Algorithm 1 Inference-only Classifier-free Guidance
@@ IV-B Setups
Following [12], both the Encodec and Watermark Encodec models use 4 RVQ codebooks, each with a vocabulary size of 2048. They operate with a stride of 320 samples, resulting in a codec framerate of 50Hz for audio recorded at a 16kHz sampling rate. The base dimension is 64, doubling at each of the 5 convolutional layers in the encoder. The number of spans to be masked, denoted as $P$, is uniformly sampled between 1 and 3. here can be intervals between different spans. The maximum masking length is set to $90\%$ of the original audio length. During training, we apply a probability of 0.5 to enhance TTS training. Text transcripts are phonemized using an IPA phoneset toolkit [^5] https://github.com/bootphon/phonemizer [27].
@@
For inference, we use nucleus sampling [29] with $p=0.8$ and a temperature of 1. The extended masked span $\alpha$ is set to $0.12$ seconds. Based on initial experiments, we determined the optimal value for the hyperparameter $\gamma$ to be $1.5$ and $\beta$ to be $5$.
@@ IV-E Results
Table I presents the results of the speech editing evaluations on RealEdit. SSR-Speech outperforms the baselines across all metrics. From the ablation studies, we observed that inference-only CFG significantly contributes to the performance improvement, effectively resolving the long silence issue in the AR model. By reducing the probabilities of undesired tokens, inference-only CFG ensures more stable and natural generation. From the MOSNet-M results, this advantage is especially pronounced in multi-span editing. Comparing with VoiceCraft that would run inference for 10 times using different margin parameters, our proposed SSR-Speech is able to inference once and obtain a stable result.
Context-aware decoding (CD) also enhances performance, particularly in speech with background sounds, as the unedited spans provide additional context for the model. Therefore, SSR-Speech with CD gets much better MOSNet-N results than the others, which bypasses the quantizer for unedited spans, preserving the unchanged parts of the audio more effectively. While watermarking (WM) does not impact performance, it adds frame-level watermarks to the synthesized audio, increasing the model’s security. Both CFG and watermark encoding independently contribute to improved performance in speech editing, while their combination achieves the best results.